OpenStudy (anonymous):

x^2 - 7x what number would i add to the polynomial to complete the sqaure? the answer i got was 12.25 is that correct?

OpenStudy (anonymous):

you would add (-7/2)^2 = (7/2)^2 = 49/4 = (48 + 1)/4 = 12.25 correct

OpenStudy (anonymous):

thank you sir.

OpenStudy (anonymous):

no problem ma'am :)